OpenStudy (anonymous):

Hooke's Law. The distance d when a spring is streched by a hanging object varies directly as weight w of the object. If the distance is 26 cm when the weight is 3 kg, what is the distance when the weight is 8kg? The distance is ? cm.

OpenStudy (anonymous):

8/3 X 26 =69.333cm

OpenStudy (amistre64):

d = kw 26 = k(3) k = 26/3 d = (26/3)(8)

OpenStudy (amistre64):

same thing

OpenStudy (anonymous):

d = kw 26 = k(3) k = 26/3 d = (26/3)(8)